# Analytic and Variational Stability of Deep Learning Systems

### v1

@@ Abstract
We propose a unified analytic and variational framework for studying stability in deep learning systems viewed as coupled representation–parameter dynamics. The central object is the Learning Stability Profile, which tracks the infinitesimal response of representations, parameters, and update mechanisms to perturbations along the learning trajectory. We prove a Fundamental Analytic Stability Theorem showing that uniform boundedness of these stability signatures is equivalent, up to norm equivalence, to the existence of a Lyapunov-type energy that dissipates along the learning flow.
In smooth regimes, the framework yields explicit stability exponents linking spectral norms, activation regularity, step sizes, and learning rates to contractivity of the learning dynamics. Classical spectral stability results for feedforward networks, a discrete CFL-type condition for residual architectures, and parametric and temporal stability laws for stochastic gradient methods arise as direct consequences. The theory extends to non-smooth learning systems, including ReLU networks, proximal and projected updates, and stochastic subgradient flows, by replacing classical derivatives with Clarke generalized derivatives and smooth energies with variational Lyapunov functionals.
The resulting framework provides a unified dynamical description of stability across architectures and optimization methods, clarifying how architectural and algorithmic choices jointly govern robustness and sensitivity to perturbations. It also provides a foundation for further extensions to continuous-time limits and geometric formulations of learning dynamics.
Keywords Analytic stability $\cdot$ learning dynamics $\cdot$ Lyapunov methods $\cdot$ energy-dissipative systems $\cdot$ generalization theory $\cdot$ deep neural networks $\cdot$ dynamical systems.
@@ 1 Introduction
Stability plays a central role in learning theory, but is formalized differently across fields. In numerical analysis, stability refers to bounded responses of discrete evolutions to perturbations and is certified by Lyapunov energies [HirschSmaleDevaney2012, Strikwerda2004]. In statistical learning theory, stability is defined through algorithmic sensitivity to data perturbations, beginning with [Bousquet2002] and extended to stochastic gradient methods by [Hardt2016Train]. A separate literature studies stability through Jacobian spectra and curvature using tools from dynamical systems and control [Sontag1998]. Recent empirical work further shows that explanations and attribution scores of modern networks can be highly unstable under perturbations of inputs, parameters, or training conditions [Fel2022, Gan2025].
Despite their maturity, these perspectives rely on distinct primitives and do not provide a unified mathematical description of stability in deep learning. This paper introduces a single analytic and variational framework that unifies these views.
The central construct is the Learning Stability Profile (LSP), which records how infinitesimal perturbations propagate through the coupled representation, parameter, and update dynamics of learning. The LSP is defined at the level of the full learning flow and applies uniformly to smooth and non-smooth architectures, including ReLU networks, proximal and projected methods, stochastic subgradient flows, residual networks, and continuous-time limits. Our main result shows that, up to equivalence of norms, boundedness of the Learning Stability Profile is equivalent to the existence of a Lyapunov-type energy that dissipates along the learning trajectory.
This equivalence yields concrete stability laws. In smooth regimes it recovers conditions based on Jacobian bounds, spectral regularity, and step-size constraints. For residual networks it produces a discrete Courant–Friedrichs–Lewy type restriction consistent with explicit time stepping [Strikwerda2004]. In stochastic and non-smooth regimes, the same analytic structure yields stability statements consistent with stochastic approximation theory [KushnerYin2003, Benaim1999] and generalized variational analysis [Clarke1990Optimization, RockafellarWets1998].
The contribution is therefore not an architecture-specific condition, but a unified analytic principle connecting perturbation growth, stability exponents, and variational energy dissipation across modern learning systems.
@@ 2 Setup
We model learning as a coupled discrete-time dynamical system with state
@@
where $X_{k,t}$ denotes representations, $\theta_{k,t}$ parameters, and $U_{k,t}$ update variables such as optimization states or stochastic noise.
The system evolves through three maps: a representation map $F_{\theta}$, a parameter update map $\mathcal{A}$, and an update dynamics map $\mathcal{U}$. A single learning step is
$X_{k+1,t}=F_{\theta_{k,t}}(X_{k,t}),\qquad\theta_{k,t+1}=\mathcal{A}(\theta_{k,t},U_{k,t}),\qquad U_{k,t+1}=\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t}).$
Derivatives are classical when the maps are smooth and understood in the Clarke sense when they are only locally Lipschitz [Clarke1990Optimization]. When updates involve randomness, the system is interpreted as a measurable stochastic process [KushnerYin2003, Benaim1999].
To quantify sensitivity, we consider perturbations of the initial state $Z_{0,0}$. The first-order response of a single learning step admits the block representation
$\mathcal{J}_{k,t}=\begin{pmatrix}D_{X}F_{\theta}&D_{\theta}F_{\theta}&0\\ 0&D_{\theta}\mathcal{A}&D_{U}\mathcal{A}\\ D_{X}\mathcal{U}&D_{\theta}\mathcal{U}&D_{U}\mathcal{U}\end{pmatrix}$,
@@
interpreted as a classical or generalized Jacobian. The sequence $(\mathcal{J}_{k,t})$ constitutes the Learning Stability Profile and encodes perturbation propagation jointly across depth and time.
Asymptotic growth of these sensitivities is summarized by analytic stability exponents. For representations,
$\alpha_{x}=\limsup_{n\to\infty}\frac{1}{n}\log\bigl\|D_{Z_{0,0}}X_{n,0}\bigr\|,$
with analogous definitions for $\alpha_{\theta}$ and $\alpha_{u}$. Nonpositive exponents correspond to stability and strictly negative exponents to strict stability [HirschSmaleDevaney2012].
A variational learning energy is a function $\mathcal{E}(Z)$ satisfying a dissipation inequality
$\mathcal{E}(Z_{k+1,t})-\mathcal{E}(Z_{k,t})\leq-\gamma\,\Psi(Z_{k,t}),$
for some $\gamma>0$ and nonnegative $\Psi$. This generalizes discrete Lyapunov theory in the variational framework of [RockafellarWets1998]. The next section formalizes the Learning Stability Profile and establishes the analytic stability theorem.
Definition 2.1 (Learning System).
A learning system is specified by maps
$F_{\theta}:\mathcal{X}_{k}\to\mathcal{X}_{k+1},\qquad\mathcal{A}:\Theta\times\mathcal{U}\to\Theta,\qquad\mathcal{U}:\mathcal{U}\times\mathcal{X}_{0}\times\Theta\to\mathcal{U},$
which induce the joint state evolution
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t})=\bigl(F_{\theta_{k,t}}(X_{k,t}),\,\mathcal{A}(\theta_{k,t},U_{k,t}),\,\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t})\bigr)$
on the product space
$\mathcal{M}=\mathcal{X}_{0}\times\cdots\times\mathcal{X}_{L}\times\Theta\times\mathcal{U}.$
A trajectory is any sequence $(Z_{k,t})$ obtained by iterating $\Phi_{k,t}$ from admissible initial conditions.
Assumption 1 (Analytic Regularity).
Each map $\Phi_{k,t}$ is either continuously differentiable or locally Lipschitz with a nonempty Clarke generalized Jacobian. Constants
$L_{\mathrm{fwd}},\qquad L_{\mathrm{par}},\qquad L_{\mathrm{upd}}$
exist such that for all $(x,\theta,u)$, and every Jacobian or generalized Jacobian element
$V\in\partial\Phi_{k,t}(x,\theta,u),$
the corresponding blocks satisfy
$\|V_{x}\|\leq L_{\mathrm{fwd}},\qquad\|V_{\theta}\|\leq L_{\mathrm{par}},\qquad\|V_{u}\|\leq L_{\mathrm{upd}}.$
This ensures uniformly bounded directional derivatives in the representation, parameter, and update directions.
3 Learning Stability Profile
Section 2 models learning as a discrete-time dynamical system on a joint state space. At layer index $k$ and training step $t$, the state is
$Z_{k,t}=(X_{k,t},\theta_{k,t},U_{k,t}),$
and one step of the learning flow is
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t})=\bigl(F_{\theta_{k,t}}(X_{k,t}),\,\mathcal{A}(\theta_{k,t},U_{k,t}),\,\mathcal{U}(U_{k,t},X_{k,t},\theta_{k,t})\bigr).$
By Assumption 1, each $\Phi_{k,t}$ is either continuously differentiable or locally Lipschitz with generalized derivatives in the sense of [Clarke1990Optimization]. This section defines the Learning Stability Profile, which quantifies the propagation of infinitesimal perturbations in the input, parameter, and update directions.
3.1 Directional sensitivities of the learning flow
Define the product manifold
$\mathcal{M}=\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\times\mathcal{U},$
equipped with any compatible norm. Since $\mathcal{M}$ is finite dimensional, all such norms are equivalent.
Fix an architecture and training protocol. Consider three perturbations: an input perturbation $\delta x$ of $X_{0,0}$, an initialization perturbation $\delta\theta^{0}$ of $\theta_{0,0}$, and an update perturbation $\delta u$ of the update sequence $(U_{k,t})$. Under Assumption 1, the dependence of $Z_{k,t}$ on these perturbations is differentiable in the classical or Clarke sense [Clarke1990Optimization].
The associated directional sensitivity operators are
$\mathsf{S}^{x}_{k,t}=\frac{\partial Z_{k,t}}{\partial x}\Big|_{(x_{0},\theta^{0},u)},\qquad\mathsf{S}^{\theta}_{k,t}=\frac{\partial Z_{k,t}}{\partial\theta^{0}}\Big|_{(x_{0},\theta^{0},u)},\qquad\mathsf{S}^{u}_{k,t}=\frac{\partial Z_{k,t}}{\partial u}\Big|_{(x_{0},\theta^{0},u)}.$
In the smooth case these coincide with classical derivatives of the composite map $(x,\theta^{0},u)\mapsto Z_{k,t}$. In the non-smooth case they are elements of the Clarke generalized Jacobian [Clarke1990Optimization, RockafellarWets1998].
The collection
$\bigl\{\mathsf{S}^{x}_{k,t},\,\mathsf{S}^{\theta}_{k,t},\,\mathsf{S}^{u}_{k,t}\bigr\}_{k,t}$
constitutes the Learning Stability Profile and encodes linearized perturbation propagation across depth and training time.
3.2 Analytic stability exponents
We now compress the local stability signatures into asymptotic exponents. These quantities measure the effective growth or decay of infinitesimal perturbations along both depth and training time, and play a role analogous to Lyapunov exponents in dynamical systems [HirschSmaleDevaney2012].
Definition 3.1 (Analytic stability exponents).
For a fixed architecture and training protocol, let $L$ and $T$ denote network depth and training horizon. The analytic forward exponent is
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\Bigr).$
The analytic parameter exponent is
$\alpha_{\theta}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{\theta}_{k,t}\Bigr),$
and the analytic update exponent is
$\alpha_{u}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\Bigl(\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{u}_{k,t}\Bigr).$
These exponents are well defined under Assumption 1, since the learning flow is locally Lipschitz and the stability signatures are finite along any finite horizon. They represent large-scale growth rates of infinitesimal perturbations. Negative values correspond to exponential decay, nonpositive values to uniform boundedness, and positive values to instability.
Analytic stability is expressed as constraints on these exponents.
Definition 3.2 (Analytic stability and stability indices).
The learning system is analytically stable if
$\alpha_{x}\leq 0,\qquad\alpha_{\theta}\leq 0,\qquad\alpha_{u}\leq 0.$
It is strictly analytically stable if all three exponents are strictly negative. A triple $(\beta_{x},\beta_{\theta},\beta_{u})$ with $\beta_{x},\beta_{\theta},\beta_{u}\leq 0$ is called a stability index if
$\alpha_{x}\leq\beta_{x},\qquad\alpha_{\theta}\leq\beta_{\theta},\qquad\alpha_{u}\leq\beta_{u}.$
Definition 3.3 (Analytic learning energy).
A functional
$\mathcal{E}:\mathcal{X}_{0}\times\dots\times\mathcal{X}_{L}\times\Theta\to\mathbb{R}_{+}$
is called an analytic learning energy if it satisfies the following conditions.
  1. 1.
    Coercivity. There exist constants $c_{1},c_{2}>0$ such that, for all admissible states,
    $c_{1}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr)\leq\mathcal{E}(X_{0},\dots,X_{L},\theta)\leq c_{2}\bigl(\|X_{L}\|^{2}+\|\theta\|^{2}\bigr).$
  2. 2.
    Differentiability. The map $\mathcal{E}$ is continuously Fréchet differentiable with respect to each of its arguments.
  3. 3.
    Compatibility with the learning flow. For every layer $k$, the directional derivatives $\nabla_{X_{k}}\mathcal{E}$ and $\nabla_{\theta}\mathcal{E}$ exist and are well defined along the learning trajectory generated by the update maps $f_{k}$ and $\mathcal{A}_{t}$.
A functional satisfying these properties couples the discrete learning dynamics to a smooth Lyapunov structure.
Assumption 2 (Analytic energy–dissipation law).
There exist an analytic learning energy $\mathcal{E}$, a constant $\gamma>0$, and a perturbation remainder $R$, linear at first order in the initial perturbations $(\delta x,\delta\theta^{0},\delta u)$, such that for every training step $t$ and every admissible trajectory,
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\bigl(\|X_{L}^{t}\|^{2}+\|\theta^{t}\|^{2}\bigr)+R\bigl(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u\bigr).$
The constant $\gamma$ is the dissipation rate. The remainder satisfies the uniform bound
$|R|\leq C\bigl(\|\delta x\|+\|\delta\theta^{0}\|+\|\delta u\|\bigr)$
for some $C>0$ independent of depth, training horizon, and state.
The following basic properties make these definitions robust.
Proposition 3.4 (Independence of norm).
Let $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ be two norms on $\mathcal{M}$ inducing operator norms $\|\cdot\|_{1,\mathrm{op}}$ and $\|\cdot\|_{2,\mathrm{op}}$, associated stability signatures $\sigma^{x,1}_{k,t}$, $\sigma^{x,2}_{k,t}$, and exponents $\alpha_{x}^{(1)}$, $\alpha_{x}^{(2)}$. Then there exist constants $c,C>0$ such that for all $k,t$
$c\,\sigma^{x,1}_{k,t}\leq\sigma^{x,2}_{k,t}\leq C\,\sigma^{x,1}_{k,t},$
and the exponents coincide, $\alpha_{x}^{(1)}=\alpha_{x}^{(2)}$. The same conclusion holds for $\alpha_{\theta}$ and $\alpha_{u}$.
@@ 3.2 Analytic stability exponents
Since $\mathcal{M}$ is finite dimensional, the norms $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ are equivalent, so there exist $c,C>0$ such that
$c\|v\|_{1}\leq\|v\|_{2}\leq C\|v\|_{1}\qquad\text{for all }v\in\mathcal{M}.$
This implies corresponding bounds on the induced operator norms and hence on the stability signatures. Taking logarithms, dividing by $L+T$, and passing to the limsup removes additive constants, yielding equality of the exponents. ∎
Proposition 3.5 (Finite horizon bound and exponent sign).
The following statements hold.
  1. 1.
    If there exists a constant $C_{x}\geq 1$ such that $\sigma^{x}_{k,t}\leq C_{x}$ for all $k,t$, then $\alpha_{x}\leq 0$. Conversely, if $\alpha_{x}<0$, then there exist $\varepsilon>0$ and $C^{\prime}_{x}\geq 1$ such that for all sufficiently large $L+T$,
    $\sup_{0\leq k\leq L,\,0\leq t\leq T}\sigma^{x}_{k,t}\leq C^{\prime}_{x}\,\mathrm{e}^{-\varepsilon(L+T)}.$
  2. 2.
    The same statements hold with $\sigma^{\theta}_{k,t}$ and $\alpha_{\theta}$, and with $\sigma^{u}_{k,t}$ and $\alpha_{u}$.
Proof.
If $\sigma^{x}_{k,t}\leq C_{x}$ uniformly, then for all $L,T$
$\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)\leq\frac{1}{L+T}\log C_{x},$
and taking the limsup gives $\alpha_{x}\leq 0$. Conversely, if $\alpha_{x}<0$, there exist $\varepsilon>0$ and $N$ such that for all $L+T\geq N$,
$\frac{1}{L+T}\log\Bigl(\sup_{k,t}\,\sigma^{x}_{k,t}\Bigr)\leq-\varepsilon.$
Rearranging yields the stated exponential bound, with $C^{\prime}_{x}$ chosen to absorb the finitely many cases $L+T<N$. The arguments for $\alpha_{\theta}$ and $\alpha_{u}$ are identical. ∎
These results show that analytic stability may be characterized either by uniform boundedness of sensitivity operators along the learning trajectory or by nonpositivity of the associated asymptotic exponents. Later sections establish that these equivalent conditions are, in turn, equivalent to the existence of Lyapunov-type energies satisfying discrete dissipation inequalities. This links the local, infinitesimal profile developed here to the global energy structures of classical stability theory [HirschSmaleDevaney2012, RockafellarWets1998].
@@ 4 Analytic Stability Theorem
We now connect the Learning Stability Profile of Section 3 with Lyapunov-type energy functionals. Throughout this section we work under Assumption 1 and use the notation
@@
with norms on product spaces understood to be equivalent Euclidean norms.
The stability exponents of Definition 3.1 summarize the worst-case growth of the three components of the Learning Stability Profile. The following theorem shows that nonpositive exponents are equivalent to the existence of a dissipative analytic learning energy in the sense of Definition 3.3 and Assumption 2. The equivalence holds up to a change of norm on $\mathcal{M}$ and requires no convexity or stationarity beyond Assumption 1.
@@
Consider a learning system in the sense of Definition 2.1. Suppose that Assumption 1 holds and that the Learning Stability Profile of Definition 3.1 is well defined. Then, up to equivalence of norms on $\mathcal{M}$, the following statements are equivalent.
@@
    There exist an analytic learning energy $\mathcal{E}$ in the sense of Definition 3.3 and a constant $\gamma>0$ such that every trajectory of the coupled learning flow satisfies the dissipation inequality of Assumption 2,
@@
    where $R$ is first order in $(\delta x,\delta\theta^{0},\delta u)$.
Moreover, when these conditions hold there exist constants $c_{1},c_{2}>0$, depending only on the regularity bounds in Assumption 1, such that the stability indices of Definition 3.2 may be chosen with
$\beta_{x}\leq-c_{1}\gamma,\qquad\beta_{\theta}\leq-c_{1}\gamma,\qquad\beta_{u}\leq-c_{2}\gamma.$
@@
We work up to equivalence of norms on the finite dimensional space $\mathcal{M}$.
Assume first that item (1) holds. The uniform derivative bounds imply that each one-step map $S^{t}\mapsto S^{t+1}$ is globally Lipschitz under a suitable product norm, with constant controlled by $(C_{x},C_{\theta},C_{u})$. The system is therefore incrementally stable in the sense of [Angeli2002Incremental] and admits a smooth Lyapunov function that decays along differences of trajectories. Standard converse Lyapunov results for discrete-time systems yield a coercive $C^{1}$ functional $\mathcal{E}$ on $\mathcal{M}$ satisfying
$\mathcal{E}(S^{t+1}_{\cdot})-\mathcal{E}(S^{t}_{\cdot})\leq-\gamma\|S^{t}_{\cdot}\|^{2}+R(S^{t}_{\cdot};\delta x,\delta\theta^{0},\delta u),$
for some $\gamma>0$ and a perturbation remainder linear at first order. This $\mathcal{E}$ is an analytic learning energy, establishing item (2).
Assume now that item (2) holds. Linearizing the learning flow and the energy $\mathcal{E}$ along an arbitrary trajectory yields a discrete-time Lyapunov inequality for the linearized system. If the differential response in any of the $(x,\theta,u)$ directions were unbounded, this inequality would be violated for sufficiently large time. A contradiction argument using standard discrete-time Lyapunov theory [Khalil2002Nonlinear, Chapter 4] yields uniform bounds on the operator norms defining $\sigma^{x}_{k,t}$, $\sigma^{\theta}_{k,t}$, and $\sigma^{u}_{k,t}$. Hence the exponents $(\alpha_{x},\alpha_{\theta},\alpha_{u})$ are nonpositive and item (1) follows. The bounds on $(\beta_{x},\beta_{\theta},\beta_{u})$ follow from norm equivalence and exponential decay induced by $\gamma$. ∎
Theorem 4.1 admits three natural projections, each isolating one coordinate of the Learning Stability Profile and its associated exponent.
Corollary 4.2 (Forward analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The forward exponent satisfies $\alpha_{x}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{x}\geq 1$ such that for every input perturbation $\delta x$, every depth $k$ and every time $t$,
    $\|\delta X^{t}_{k}\|\leq C_{x}\|\delta x\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta x$ at first order.
Corollary 4.3 (Parametric analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The parametric exponent satisfies $\alpha_{\theta}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{\theta}\geq 1$ such that for every initialization perturbation $\delta\theta^{0}$, every depth $k$ and every time $t$,
    $\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{\theta}\|\delta\theta^{0}\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta\theta^{0}$ at first order.
Corollary 4.4 (Temporal analytic stability law).
Under the hypotheses of Theorem 4.1 the following statements are equivalent.
  1. 1.
    The temporal exponent satisfies $\alpha_{u}\leq 0$.
  2. 2.
    There exist a learning energy $\mathcal{E}$ and a constant $C_{u}\geq 1$ such that for every admissible perturbation of the update mechanism $\delta u$, and for all depths $k$ and times $t$,
    $\|\delta X^{t}_{k}\|+\|\delta\theta^{t}\|\leq C_{u}\|\delta u\|,$
    and the dissipation inequality of Assumption 2 holds with a perturbation remainder that is linear in $\delta u$ at first order.
These corollaries show that forward robustness, parametric robustness, and temporal robustness are not independent notions. Each is a coordinate projection of the same analytic object, namely the Learning Stability Profile, and each is equivalent to the existence of a Lyapunov-type energy that dissipates along the corresponding perturbation direction.

### v2

# Analytic and Variational Stability in Deep Learning Systems

@@ Abstract
We propose a unified analytic and variational framework for stability in deep learning systems viewed as coupled representation–parameter dynamics. The central object is the Learning Stability Profile, which measures how infinitesimal perturbations propagate through representations, parameters, and update mechanisms along the learning trajectory.
Our main result, the Fundamental Analytic Stability Theorem, shows that uniform boundedness of these sensitivities is equivalent, up to norm equivalence, to the existence of a Lyapunov-type energy dissipating along the learning flow. In smooth regimes, this yields explicit stability exponents linking spectral norms, activation regularity, step sizes, and learning rates to contractive behavior. Classical spectral stability of feedforward networks, CFL-type conditions for residual architectures, and temporal stability laws for stochastic gradient methods follow as direct consequences.
The framework extends to non-smooth systems, including ReLU networks, proximal and projected updates, and stochastic subgradient flows, by replacing classical derivatives with Clarke generalized derivatives and smooth energies with variational Lyapunov functionals.
The resulting theory provides a unified dynamical description of stability across architectures and optimization methods, clarifying how design and training choices jointly control robustness and sensitivity to perturbations.
Keywords learning dynamics $\cdot$ stability analysis $\cdot$ Lyapunov methods $\cdot$ deep neural networks $\cdot$ stochastic gradient methods.
@@ 1 Introduction
Stability is a central concept in learning theory, but is formalized differently across numerical analysis, statistical learning, and dynamical systems. Classical notions include Lyapunov stability for discrete and continuous evolutions Hirsch et al. [2012]; Strikwerda [2004], algorithmic stability under data perturbations Bousquet and Elisseeff [2002]; Hardt et al. [2016], and Jacobian- or curvature-based criteria from control theory Sontag [1998]. These approaches address different objects and do not yield a unified analytic description of stability for modern learning systems.
We introduce a single framework that connects these notions. The central object is the Learning Stability Profile (LSP), which captures how infinitesimal perturbations propagate along the full learning flow, jointly over representations, parameters, and update dynamics. The LSP is architecture-independent and applies to smooth, nonsmooth, and stochastic learning systems, and admits continuous-time limits.
Our main result establishes an equivalence, up to norm equivalence, between boundedness of the Learning Stability Profile and the existence of a Lyapunov-type energy that dissipates along the learning trajectory. This yields a necessary and sufficient characterization of stability at the level of the learning dynamics itself. Classical stability conditions in smooth, residual, stochastic, and variational regimes are recovered as special cases Strikwerda [2004]; Kushner and Yin [2003]; Benaïm [1999]; Clarke [1990]; Rockafellar and Wets [1998].
@@ 2 Setup
We model learning as a discrete-time dynamical system on a joint state space
@@
with evolution
$Z_{k+1,t+1}=\Phi_{k,t}(Z_{k,t}).$
The maps $\Phi_{k,t}$ are assumed to be either smooth or locally Lipschitz, with generalized derivatives interpreted in the sense of Clarke when necessary, and stochastic interpretations when applicable Clarke [1990]; Kushner and Yin [2003]; Benaïm [1999].
@@
Let $\mathcal{J}_{k,t}$ denote a (classical or generalized) Jacobian of $\Phi_{k,t}$. The sequence $(\mathcal{J}_{k,t})$ describes linearized perturbation propagation jointly across depth and training time and constitutes the Learning Stability Profile.
We quantify long-horizon sensitivity growth through analytic stability exponents. For example, the forward exponent is
$\alpha_{x}=\limsup_{L,T\to\infty}\frac{1}{L+T}\log\sup_{0\leq k\leq L,\;0\leq t\leq T}\bigl\|\partial_{Z_{0,0}}X_{k,t}\bigr\|,$
with analogous definitions for parameter and update directions. Nonpositive exponents correspond to uniform boundedness, and negative exponents to exponential decay Hirsch et al. [2012].
We also consider Lyapunov-type energies compatible with the learning flow. An analytic learning energy is a coercive, differentiable functional whose variation along trajectories satisfies a discrete dissipation inequality, in the sense of variational stability theory Rockafellar and Wets [1998].
Proposition 2.1 (Well-posedness).
The analytic stability exponents are independent of the choice of norm on the finite-dimensional state space. Uniform boundedness of linearized sensitivities is equivalent to nonpositivity of the corresponding exponents, and strict negativity implies exponential decay along sufficiently long depth–time horizons.
@@
All statements follow directly from equivalence of norms in finite dimensions and from the definition of the exponents as normalized logarithmic growth rates. ∎
The next section formalizes the Learning Stability Profile and establishes the equivalence between bounded analytic stability exponents and the existence of dissipative learning energies.
@@ 3 Analytic Stability Theorem
We now connect the Learning Stability Profile with Lyapunov-type energy functionals. Throughout this section we work with a learning system whose update maps are either smooth or locally Lipschitz, with generalized derivatives interpreted in the usual variational sense. We use the notation
@@
and work up to equivalence of norms on this finite-dimensional product space.
The stability exponents summarize worst-case growth of infinitesimal perturbations in the representation, parameter, and update directions. The following theorem shows that nonpositive exponents are equivalent to the existence of a dissipative learning energy. The equivalence holds up to a change of norm and requires no convexity, stationarity, or algorithm-specific structure.
@@
Consider a learning system whose Learning Stability Profile is well defined. Then, up to equivalence of norms on $\mathcal{M}$, the following statements are equivalent.
@@
    There exist a coercive, differentiable learning energy $\mathcal{E}$ and a constant $\gamma>0$ such that every trajectory of the learning flow satisfies
@@
    where $R$ is linear at first order in $(\delta x,\delta\theta^{0},\delta u)$.
Moreover, when these conditions hold there exist constants $c_{1},c_{2}>0$, depending only on the regularity bounds of the learning flow, such that the stability exponents satisfy
$\alpha_{x}\leq-c_{1}\gamma,\qquad\alpha_{\theta}\leq-c_{1}\gamma,\qquad\alpha_{u}\leq-c_{2}\gamma.$
@@
All arguments are carried out up to equivalence of norms on $\mathcal{M}$.
Assume first that uniform bounds on the sensitivity operators hold. Then the one-step learning map is globally Lipschitz under a suitable product norm. This implies incremental stability of the learning dynamics and, by standard converse Lyapunov results for discrete-time systems, yields a coercive $C^{1}$ energy functional whose value decreases along trajectories, up to a first-order perturbation remainder. This establishes the existence of a dissipative learning energy.
Conversely, suppose such an energy exists. Linearizing the learning flow and the energy along an arbitrary trajectory yields a discrete Lyapunov inequality for the linearized system. Unbounded growth of perturbations in any coordinate direction would violate this inequality for sufficiently large horizons. A contradiction argument based on classical discrete-time Lyapunov theory implies uniform bounds on all sensitivity operators, and hence nonpositivity of the stability exponents. The quantitative bounds on the stability indices follow from norm equivalence and the dissipation rate $\gamma$. ∎
Theorem 3.1 admits natural projections isolating individual coordinates of the Learning Stability Profile and their associated exponents.